OpenStudy (anonymous):

Please help!!! A dealer has a maximum of $4000 allocated to spend on the purchase of two types of car accessories. Minimum orders from a warehouse is 25 of type A and 20 of Type B. Discounted price for purchase is $25 per unit of type A and $40 per unit of Type B. They can be retailed at a profit of $20 per unit of Type A and $35 per unit of Type B. How many of each type should he buy to maximise profits? I'm really stuck on this please show me what to do :( anyone.

OpenStudy (anonymous):

ok thanks if not no biggie... gonna head to bed goodnight u guys...P:)

OpenStudy (anonymous):

It does look confusing. I'll see if I can decipher it and comment on it. :-) G'nite!! :-)

OpenStudy (anonymous):

:) ciao

OpenStudy (anonymous):

Okay, it looks like the solution can be laid out like this: (4000/type_cost) is how many he can purchase of a given type. E.g.: for type-A, he can purchase (4000/25) = 160 units. This also means that: gross_yield = (4000/type_cost)*retail_markup yield_A = (4000/25)*45 yield_B = (4000/40)*75 Based on ONLY the information you've supplied, I see no way that any combination of purchased parts could yield more than blowing the entire 4 grand on "Type B" units.